# Driving of an open quantum system at finite temperature across first- and second-order quantum phase transitions

### v2

@@ Abstract
An open fully connected system of qubits at nonzero temperature is driven within a finite time interval along various paths in the space of its control parameters. The driving leads across finite-size precursors of the first- and second-order quantum phase transition from factorized to entangled ground-state phases, aiming at the preparation of the complex ground state of the system at the final parameter point with maximal fidelity. During the drive, the system is coupled to a heat bath with a constant temperature, the dynamics being determined in a nonpertubative way by the method of Hierarchical Equations of Motion. It is shown that the presence of the heat bath in combination with specific patterns of avoided crossings affecting the ground and excited states in the parameter region around the quantum phase transition may considerably improve the fidelity of preparation of the target ground state.
@@ 1 INTRODUCTION
In the last decades different strategies have been proposed to accomplish this task. One can mention, among the most striking approaches, the adiabatic, transitionless, geometric, and decoherence-assisted types of driving. The adiabatic driving represents a direct application of the quantum adiabatic theorem, which states that for slow enough variation of control parameters the system follows the instantaneous (adiabatic) ground state in the parameter space Kato50 ; Jans07 ; Elga12 . The transitionless (also called counter-diabatic) driving emulates the adiabatic evolution along a given parameter path with the aid of additional terms in the Hamiltonian (which usually are non-local) that compensate non-adiabatic effects of the finite-speed driving Demi03 ; Berr09 ; Camp13 ; Guer19 . The geometric driving makes use of the geodesic path between the initial and final parameter points with respect to the Provost-Vallee metric on the ground-state manifold in attempt to maximize the overlap of the evolving state with the adiabatic ground state Prov80 ; Anan90 ; Miya01 ; Reza09 ; Kolo17 ; Buko19 . Finally, the decoherence-assisted driving increases the fidelity of the final-state preparation by repeated measurement-like interactions of the driven system with a suitable ancilla Chil02 ; Roa06 ; Haco18 ; Cejn23 .
@@
Open systems are most commonly studied within the Lindblad formalism, i.e., assuming that the system is weakly coupled to a Markovian reservoir and considering the Born and rotating wave approximations Weis12 ; Alba12 ; Yama17 ; Dann18 . Although the equations obtained in this way are numerically simple to solve, their validity is subject to too stringent conditions and the results are not guaranteed to be accurate Iles14 ; Iles16 ; Abbo20 . In this work, we use the method of Hierarchical Equations of Motion (HEOM) Tani89 ; Tani06 ; Fruch16 ; Tani20 ; Lamb23 , which is based on a numerically exact solution of quantum dynamics through the influence functional formalism Feyn63 ; Cald81 ; Cald83 . This approach is applicable to a large span of driving times and allows for narrow energy gaps. The method has already been used to study quantum dissipative dynamics under time-dependent driving fields Xu09 , entanglement dynamics Ma12 as well as more complex dynamics Ishi09 ; Chen09 ; Stru11 ; Lamb19 .
@@ 2.1 Qubit system
The QPTs between these phases are of the first order for ${\chi\neq 0}$ and of the second order for ${\chi=0}$ Ribe07 . In the first-order QPT, the order parameters change discontinuously, while the ground-state energy $E_{0}(\mathbb{\Lambda})$ shows a discontinuity of the first derivative and the ground-state energy gap $\Delta_{10}(\mathbb{\Lambda})={E_{1}(\mathbb{\Lambda})-E_{0}(\mathbb{\Lambda})}$ vanishes exponentially with ${N\to\infty}$. In contrast, the second-order QPT comes through a discontinuity of the first derivative of the order parameter, which is connected with a discontinuity of the second derivative of the ground-state energy and a polynomially vanishing ground-state energy gap. In the second-order QPT, the parity of the ground state is spontaneously broken, which means that for ${\lambda>\lambda_{\rm c}(0)}$ and ${\chi=0}$ the ${N\to\infty}$ ground state becomes a degenerate parity doublet. Finite-$N$ precursors of the QPTs of the first and second order are observed in the evolution of the lowest energy levels in panels (a) and (b) of Fig. 1. This figure depicts lower parts of energy spectra of our LMG Hamiltonian for two particular paths in the ${\lambda\times\chi}$ plane, which will be used (although with a different value of $N$) in the driving protocols to be introduced in Sec. 3.
Figure 1: Energy spectra of the LMG Hamiltonian (3) for two cuts through the ${\lambda\times\chi}$ parameter plane: (a) ${\lambda\in[0,0.5]}$, ${\chi=4.8\lambda}$ and (b) ${\lambda\in[0,4]}$, ${\chi=0}$. We set ${N=40}$. The vertical line segments demarcate ${N\to\infty}$ critical points $\lambda_{\rm c}(\chi)$ of (a) the first-order QPT and (b) the second-order QPT. In panel (b) the positive and negative parity levels are drawn by full and dashed curves, respectively, visualizing the spontaneous parity breaking at the second-order QPT. All level crossings are avoided. Their patterns around both ground-state QPTs represent precursors of ESQPTs.
The ESQPT nonanalyticities, affecting in the ${N\to\infty}$ limit higher energy levels, are also present in the spectrum of Hamiltonian (3). Their finite-$N$ precursors are also seen in Fig. 1. Associated with the first-order QPT, see Fig. 1(a), two chains of avoided level crossings of excited states issue from the ${\lambda\approx\lambda_{\rm c}(\chi)}$ avoided crossing of the ground state to both ${\lambda<\lambda_{\rm c}}$ and ${\lambda>\lambda_{\rm c}}$ directions. In addition, many other avoided crossings appear in a finite region above the ESQPTs. The second-order QPT in Fig. 1(b) is accompanied by a single chain of avoided crossings in the direction ${\lambda\gtrsim\lambda_{\rm c}(0)}$. These structures, which are generic accompaniments of the first- and second-order QPTs Cejn08 ; Cejn21 , play an important role in driving-induced dynamics of the system at nonzero temperature.
@@ 2.2 Environment and qubit-environment coupling
$\hat{H}_{\text{I}}=\hat{Q}\otimes\underbrace{\sum_{k}g_{k}\big{(}\hat{b}_{k}^{\dagger}+\hat{b}_{k}\big{)}}_{\hat{X}},$ (7)
where $\hat{X}$ is a weighted position of the oscillators, with $g_{k}$ denoting non-negative coupling strengths to individual modes, and $\hat{Q}$ is the system-environment coupling operator acting in the space of qubits. The operator $\hat{Q}$ is naturally taken as a linear combination of the quasispin components $\hat{J}_{\alpha}$. The usual choice in the literature is ${\hat{Q}=\hat{J}_{z}}$ Weis12 ; Lamb19 , but with this coupling the total Hamiltonian (1) would conserve the parity (4) for ${\chi=0}$. To deal with the situations in which the qubit-environment interaction does not conserve symmetries of the qubit system (non-fundamental symmetries shall be violated in generic cases), we consider two choices:
$\hat{Q}=\hat{J}_{z}\quad\text{or}\quad\hat{Q}=\hat{J}_{x}.$ (8)
The first choice represents a modulation of energies of the qubit states while the second one induces transitions between these states.
The fluctuation and dissipation processes involved in thermalization of the system due to its interaction with the environment can be characterized by the correlation function
$C(t)$ $=$ $\text{tr}\left[\hat{X}(t)\hat{X}(0)\,\hat{\rho}_{\rm E}^{(\beta)}\right]$
@@
where $\hat{\rho}_{\rm E}^{(\beta)}$ is the thermal density operator of the environment, and ${\hat{X}(t)=e^{+i\hat{H}_{\text{E}}t}\hat{X}e^{-i\hat{H}_{\text{E}}t}}$ is the time-dependent environment operator in the interaction Hamiltonian (7). The correlation function is fully determined by
@@
which is the density of bath mode energies weighted by squares of the respective S-E coupling strengths. In the limit of a continuous distribution of bath mode energies, $J(\omega)$ becomes a smooth function of $\omega$.
It is assumed that after a certain time the correlation decays as ${\operatorname{Im}\,C(t)\sim e^{-\gamma t}}$, where $\gamma$ represents the width of the bath mode distribution (in short, the bandwidth) that determines the dissipation time scale ${\tau_{\text{E}}\equiv\gamma^{-1}}$. By imposing this condition one finds (10) in the Drude-Lorentz form Tani89 ; Ma12 ; Lamb23 :
@@
plays the role of an effective system-environment coupling strength. The use of Eq. (12) for the numerical calculations is going to be explained in the next section.
@@ 3.1 Initial state and overall evolution
Here, $\hat{H}_{\text{S}}(0)$ is the system Hamiltonian at $t=0$, and $\hat{\rho}_{\text{S}}(0)$ and $\hat{\rho}_{\text{E}}(0)$ are thermal density operators of S and E, respectively. This corresponds to the situation when in times ${t<0}$ both S and E have thermalized separately through interactions with another thermal reservoir but with no interaction between each other.
The interaction between S and E is initiated at ${t=0}$, indicating the start of the driving procedure on S, see Sec. 3.2. One can assume that the environment E is connected with the instrumentation needed to perform the drive. Since then, the S-E interaction completely dominates over the interaction with the other reservoir so that the latter can be neglected. The evolved state $\hat{\rho}(t)$ of the composite S-E system for ${t>0}$ satisfies the quantum Liouville equation
@@ 3.2 Driving protocols
Figure 2: Top panel: Relevant part of the parameter space of the qubit Hamiltonian $\hat{H}_{\rm S}(\mathbb{\Lambda})$. The dashed curve represents the ${N\rightarrow\infty}$ QPT separatrix between the ground-state phases I and II. The tilted (solid red) line segment represents the driving path across the first-order QPT. The horizontal (solid blue) line segment represents the driving path across the second-order QPT. The gray background encodes the energy gap $\Delta_{10}(\mathbb{\Lambda})$. Lower panels: The planar speed $u$ (solid line) and geometric speed $v$ (dotted line) for driving protocols A and B along both paths. The upper row (with red curves) corresponds to the path across the first-order QPT (overflow parts of the dependencies displayed in the insets), the lower row (blue curves) corresponds to the path across the second-order QPT. We set ${N=10}$ and ${t_{\rm F}=1}$.
@@
Here, $g_{\mu\nu}(\mathbb{\Lambda})$ is the metric tensor defined on the manifold of ground states $|E_{0}(\mathbb{\Lambda})\rangle$ of Hamiltonian $\hat{H}_{\rm S}(\mathbb{\Lambda})$. Assuming a non-degenerate spectrum for finite $N$, we can write
@@
where for brevity we skip all dependencies on $\mathbb{\Lambda}$. The geometric speed (18) can be understood as the speed in a curved space whose metric tensor is determined through the distance element ${dl^{2}=}{1-\left|\langle E_{0}(\mathbb{\Lambda})|E_{0}(\mathbb{\Lambda}\!+\!d\mathbb{\Lambda})\rangle\right|^{2}}$ expressing the variation of the ground-state wave function with control parameters Prov80 . While the geometric speed (18) is constant, equal to ${v=\int_{\mathbb{\Lambda}_{\rm I}}^{\mathbb{\Lambda}_{\rm F}}\sqrt{g_{\mu\nu}d\Lambda^{\mu}d\Lambda^{\mu}}/t_{\rm F}}$, the planar speed (17) varies in time, taking smaller values in the parameter domains where $g_{\mu\nu}(\mathbb{\Lambda})$ is large, i.e., particularly in regions with a small energy gap $\Delta_{10}(\mathbb{\Lambda})$. For drivings of isolated systems initiated in the ground state $|E_{0}(\mathbb{\Lambda}_{\rm I})\rangle$, the $v$=const. strategy improves the final fidelity for the target state $|E_{0}(\mathbb{\Lambda}_{\rm F})\rangle$ Matu23a ; Matu23b .
@@
In Fig. 2 we plot the two driving paths in the plane $(\lambda,\chi)$ and the speeds (17) and (18) along both paths for driving protocols A and B for the system with ${N=10}$ qubits. We note the reduction of the planar speed $u(t)$ in both first- and second-order QPT regions (with reduced values of the ground-state energy gap $\Delta_{10}$) for the $v$=const. driving procedure B.

### v4

@@ Abstract
An open fully connected system of qubits at nonzero temperature is driven within a finite time interval along various paths in the space of its control parameters. The driving leads across finite-size precursors of first- and second-order quantum phase transition from factorized to entangled ground-state phases, aiming at the preparation of the complex ground state of the system at the final parameter point with maximal fidelity. During the drive, the system is coupled to a heat bath at a constant temperature, the dynamics being determined in a nonperturbative way by the method of Hierarchical Equations of Motion. It is shown that the presence of the heat bath in combination with specific patterns of avoided crossings affecting the ground and excited states in the parameter region around the quantum phase transition may considerably improve the fidelity of preparation of the target ground state.
@@ 1 INTRODUCTION
In the last decades different strategies have been proposed to accomplish this task. One can mention, among the most striking approaches, the adiabatic, transitionless, geometric, and decoherence-assisted types of driving. The adiabatic driving represents a direct application of the quantum adiabatic theorem, which states that for a sufficiently slow variation of control parameters, the system follows the instantaneous (adiabatic) ground state in the parameter space Kato50 ; Jans07 ; Elga12 . The transitionless (also called counter-diabatic) driving emulates the adiabatic evolution along a given parameter path with the aid of additional terms in the Hamiltonian (which usually are non-local) that compensate non-adiabatic effects of the finite-speed driving Demi03 ; Berr09 ; Camp13 ; Guer19 . The geometric driving makes use of the geodesic path between the initial and final parameter points with respect to the Provost-Vallee metric on the ground-state manifold in attempt to maximize the overlap of the evolving state with the adiabatic ground state Prov80 ; Anan90 ; Miya01 ; Reza09 ; Kolo17 ; Buko19 . Finally, the decoherence-assisted driving increases the fidelity of the final-state preparation by repeated measurement-like interactions of the driven system with a suitable ancilla Chil02 ; Roa06 ; Haco18 ; Cejn23 .
@@
Open systems Breu02 ; Weis12 ; Schl07 are most commonly studied within the Lindblad formalism, i.e., assuming that the system is weakly coupled to a Markovian reservoir and considering the Born and rotating wave approximations Alba12 ; Yama17 ; Dann18 . Although the equations obtained in this way are numerically simple to solve, their validity is subject to too stringent conditions and the results are not guaranteed to be accurate Ishi05 ; Dijk12 ; Iles14 ; Iles16 ; Abbo20 ; Anto23_1 ; Anto23_2 . In this work, we use the method of Hierarchical Equations of Motion (HEOM) Tani89 ; Tani06 ; Fruch16 ; Tani20 ; Lamb23 , which is based on a numerically exact solution of quantum dynamics through the influence functional formalism Feyn63 ; Cald81 ; Cald83 . This approach is applicable to a large span of driving times and allows for narrow energy gaps. The method has already been used to study quantum dissipative dynamics under time-dependent driving fields Xu09 , entanglement dynamics Ma12 as well as more complex dynamics Ishi09 ; Chen09 ; Stru11 ; Lamb19 .
@@ 2.1 Qubit system
The QPTs between these phases are of the first order for ${\chi\neq 0}$ and of the second order for ${\chi=0}$ Ribe07 . In the first-order QPT, the order parameters change discontinuously, while the ground-state energy $E_{0}(\mathbb{\Lambda})$ shows a discontinuity of the first derivative and the ground-state energy gap $\Delta_{10}(\mathbb{\Lambda})={E_{1}(\mathbb{\Lambda})-E_{0}(\mathbb{\Lambda})}$ vanishes exponentially with ${N\to\infty}$. In contrast, the second-order QPT comes through a discontinuity of the first derivative of the order parameter, which is connected with a discontinuity of the second derivative of the ground-state energy and a polynomially vanishing ground-state energy gap. In the second-order QPT, the parity of the ground state is spontaneously broken, which means that for ${\lambda>\lambda_{\rm c}(0)}$ and ${\chi=0}$ the ${N\to\infty}$ ground state becomes a degenerate parity doublet. A more elaborate analysis of the QPTs and the ground-state geometry of Hamiltonian (3) can be found in Matu23a ; Strel24 . Finite-$N$ precursors of the QPTs of the first and second order are observed in the evolution of the lowest energy levels in panels (a) and (b) of Fig. 1. This figure depicts lower parts of energy spectra of our LMG Hamiltonian for two particular paths in the ${\lambda\times\chi}$ plane, which will be used (although with a different value of $N$) in the driving protocols to be introduced in Sec. 3.
Figure 1: Energy spectra of the LMG Hamiltonian (3) for two cuts through the ${\lambda\times\chi}$ parameter plane: (a) ${\lambda\in[0,0.5]}$, ${\chi=4.8\lambda}$ and (b) ${\lambda\in[0,4]}$, ${\chi=0}$. We set ${N=40}$. The vertical dotted line segments demarcate ${N\to\infty}$ critical points $\lambda_{\rm c}(\chi)$ of (a) the first-order QPT and (b) the second-order QPT. In panel (b) the positive and negative parity levels are drawn by full and dashed curves, respectively, visualizing the spontaneous parity breaking at the second-order QPT. All level crossings are avoided. Their patterns (highlighted by the background color) around both ground-state QPTs represent precursors of ESQPTs.
The ESQPT nonanalyticities, affecting in the ${N\to\infty}$ limit higher energy levels, are also present in the spectrum of Hamiltonian (3). Their finite-$N$ precursors are also shown in Fig. 1 (see the parts of the spectra with the colored background). Associated with the first-order QPT, see Fig. 1(a), two chains of avoided level crossings of excited states issue from the ${\lambda\approx\lambda_{\rm c}(\chi)}$ avoided crossing of the ground state to both ${\lambda<\lambda_{\rm c}}$ and ${\lambda>\lambda_{\rm c}}$ directions. In addition, many other avoided crossings appear in a finite region above the ESQPTs. The second-order QPT in Fig. 1(b) is accompanied by a single chain of avoided crossings in the direction ${\lambda\gtrsim\lambda_{\rm c}(0)}$. These structures, which are generic accompaniments of the first- and second-order QPTs Cejn08 ; Cejn21 , play an important role in driving-induced dynamics of the system at nonzero temperature.
@@ 2.2 Environment and qubit-environment coupling
$\hat{H}_{\text{I}}=\frac{1}{\sqrt{N}}\ \hat{Q}\otimes\underbrace{\sum_{k}g_{k}\big{(}\hat{b}_{k}^{\dagger}+\hat{b}_{k}\big{)}}_{\hat{X}},$ (7)
where $\hat{X}$ is a weighted position of the oscillators, with $g_{k}$ denoting non-negative coupling strengths to individual modes, and $\hat{Q}$ is the system-environment coupling operator acting in the space of qubits. The operator $\hat{Q}$ is naturally taken as a linear combination of the quasispin components $\hat{J}_{\alpha}$, for instance
$\hat{Q}=\sin\theta\,\hat{J}_{x}+\cos\theta\,\hat{J}_{z}.$ (8)
We use two choices of the angle $\theta$, namely ${\theta=0}$ and ${\theta=\frac{\pi}{2}}$. While in the former case the environment just modulates energies of the unperturbed qubit states, in the latter case it induces transitions between these states. Note that with ${\theta=0}$ and ${\chi=0}$ the total Hamiltonian (1) conserves the parity (4), so for ${\chi=0}$ the choice ${\theta=\frac{\pi}{2}}$ captures the generic situation in which the qubit-environment interaction violates non-fundamental symmetries of the qubit system. The factor in front of the expression (7) ensures proper scaling of the S-E interaction energy with an increasing number of qubits $N$.
The total Hamiltonian (1) with individual terms fixed by Eqs. (3), (6) and (7) suffers from the drawback that for infinite sizes of both the qubit and environment subsystems and for sufficiently large coupling strengths $g_{k}$ its ground state energy may cease having a lower bound Rzaz75 . In case of a finite $N$, the lower bound exists, but the whole system with increasing $g_{k}$’s may exhibit a sharp crossover from a normal phase, characterized by low average occupancies $\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle$ of all bath modes, to a “superradiant” phase, in which these occupancies get macroscopically large Hepp73 ; Wang73 ; Emary03 . This transition is experimentally verified in various laboratory realizations of the Dicke-type Hamiltonians in which the bilinear coupling of the form (7) is also present, see, e.g., Baum10 ; Nata10 .
To avoid the ground-state instability and the transition to the superradiant-like phase, the system Hamiltonian (3) is often renormalized by adding an extra term, which counterbalances the influence of the environment Cald81 ; Cald83 . This leads to a tranformation
$\hat{H}_{\rm S}(\mathbb{\Lambda})\longmapsto\hat{H}^{\prime}_{\rm S}(\mathbb{\Lambda},q)=\hat{H}_{\rm S}(\mathbb{\Lambda})+r\frac{q}{N}\hat{Q}^{2},$ (9)
where variable $r$ will be explained below and
$q=\sum_{k}\frac{g_{k}^{2}}{\omega_{k}}$ (10)
plays the role of an effective S-E coupling strength. With the replacement (9), the total Hamiltonian (1) is guaranteed to have a lower bound of energy Kopyl19 .
In this work, the qubit system is always considered finite, so the inclusion of the counterterm is not necessary. Nevertheless, as we wish to estimate the effect of renormalization in the driving process, we include the counterterm to some of our simulations. Whether the renormalization is on or off, respectively, is distinguished by values ${r=1}$ or ${r=0}$ of the auxiliary parameter in Eq. (9). In particular, the option ${r=1}$ is used along with ${\hat{Q}=\hat{J}_{x}}$, when the coupling term (7) can be interpreted as the electric dipole interaction of the qubits (atoms carrying the qubit states) with $\omega_{k}$ modes of a quantized electromagnetic field and the counterterm stands for the $A^{2}$ term (squared vector potential) of the atom-field Hamiltonian Rzaz75 ; Nata10 . Hence for ${\theta=\frac{\pi}{2}}$ in Eq. (8) we use both ${r=0}$ and ${r=1}$ options, while for ${\theta=0}$ we use only ${r=0}$. With the choice ${r=0}$, we intentionally allow for the situation in which the increasing effective strength (10) of the S-E interaction induces a crossover effect in the ground state of the whole system to a superradiant-like phase, which may have a large impact on results of the driving process.
Thermal fluctuations of the environment are described by its self-correlation function
$C(t)$ $=\text{tr}\left[\hat{X}(t)\hat{X}(0)\,\hat{\rho}_{\rm E}^{(\beta)}\right]$ (11)
$=\sum_{k}g_{k}^{2}\biggl{[}\bigl{\{}2n_{\rm E}(\omega_{k})\!+\!1\bigr{\}}\cos(\omega_{k}t)-i\sin(\omega_{k}t)\biggr{]}$
@@
where $\hat{\rho}_{\rm E}^{(\beta)}$ is a thermal density operator of the environment and ${\hat{X}(t)=e^{+i\hat{H}_{\text{E}}t}\hat{X}e^{-i\hat{H}_{\text{E}}t}}$ represents time evolution of the weighted oscillator coordinate from Eq. (7). In the second line we use the Bose distribution $n_{\rm E}(\omega_{k})=(e^{\beta\omega_{k}}-1)^{-1}$ and in the third line we introduce the spectral density
@@
which fully determines the correlation function Breu02 ; Weis12 . For discrete modes, the spectral density consists of a sequence of $\delta$-peaks, but in the limit of a continuous distribution of bath mode energies (infinite number of bosonic modes) it becomes a smooth function of $\omega$. While the real part of the correlation function (11) describes thermal fluctuations and decoherence, the imaginary part captures the effects of dissipation. For more detailed explanation we suggest Refs. Breu02 ; Tani89 ; Weis12 ; Schl07 .
It is assumed that the dissipative effect of the bath decays as ${\operatorname{Im}\,C(t)\sim e^{-\gamma t}}$, where ${\gamma^{-1}=\tau_{\text{E}}}$ is the dissipation time scale. By imposing this condition one finds (12) in the Drude-Lorentz form Tani89 ; Ma12 ; Lamb23 :
@@
The use of this formalism in numerical calculations based on the HEOM method will be further commented in Sec. 3.4.
@@ 3.1 Initial state and overall evolution
Here, $\hat{H}_{\text{S}}(0)$ is the system Hamiltonian at $t=0$, and $\hat{\rho}_{\text{S}}(0)$ and $\hat{\rho}_{\text{E}}(0)$ are thermal density operators of S and E, respectively. This corresponds to the situation when in times ${t<0}$ both S and E have thermalized separately through interactions with a common thermal reservoir but with no interaction between each other. We note that the thermal initial state of S in Eq. (14) is probably a more realistic choice than a pure zero-temperature state ${\hat{\rho}_{\rm S}(0)=|E_{0}(0)\rangle\langle E_{0}(0)|}$. Moreover, it will be shown that the nonzero temperature of S has a non-trivial effect on the final fidelity of the driving procedure.
The interaction between S and E is initiated at ${t=0}$, indicating the start of the driving procedure on S, see Sec. 3.2. One can assume that the environment E is connected with the instrumentation needed to perform the drive. Since then, the S-E interaction completely dominates over the interaction with the other reservoir so that the latter can be neglected. The evolved state $\hat{\rho}(t)$ of the composite S-E system for ${t>0}$ satisfies the quantum Liouville–von Neumann equation
@@ 3.2 Driving protocols
Figure 2: Top panel: Relevant part of the parameter space of the qubit Hamiltonian $\hat{H}_{\rm S}(\mathbb{\Lambda})$. The dashed curve represents the ${N\rightarrow\infty}$ QPT separatrix between the ground-state phases I and II. The tilted (solid red) line segment represents the driving path across the first-order QPT. The horizontal (solid blue) line segment represents the driving path across the second-order QPT. The gray background encodes the energy gap $\Delta_{10}(\mathbb{\Lambda})$ for ${N=10}$. Lower panels: The planar speed $u$ (solid line) and geometric speed $v$ (dotted line) for driving protocols A and B along both paths for ${N=10}$ and ${t_{\rm F}=1}$. The upper and lower rows (with red and blue curves) correspond to the paths across the first- and second-order QPTs, respectively (overflow parts of the dependencies are displayed in the insets). The vertical dashed line segments mark the crossing of the ${N\rightarrow\infty}$ QPT separatrix from the upper panel.
@@
Here, $g_{\mu\nu}(\mathbb{\Lambda})$ is the metric tensor defined on the manifold of ground states $|E_{0}(\mathbb{\Lambda})\rangle$ of Hamiltonian $\hat{H}_{\rm S}(\mathbb{\Lambda})$ via the distance element ${dl^{2}=1-\left|\langle E_{0}(\mathbb{\Lambda})|E_{0}(\mathbb{\Lambda}\!+\!d\mathbb{\Lambda})\rangle\right|^{2}}$ Prov80 . Assuming a non-degenerate spectrum for finite $N$, we can write
@@
where for brevity we skip all dependencies on $\mathbb{\Lambda}$.
The metric (20) and the geometric speed (19) reflect variations of the ground-state wave function with control parameters $\mathbb{\Lambda}$. In protocol B, the geometric speed is kept constant, equal to $v=\int_{\mathbb{\Lambda}_{\rm I}}^{\mathbb{\Lambda}_{\rm F}}\sqrt{g_{\mu\nu}d\Lambda^{\mu}d\Lambda^{\mu}}/t_{\rm F}$, while the planar speed (18) varies in time, taking smaller values in the parameter domains where $g_{\mu\nu}(\mathbb{\Lambda})$ is large, i.e., particularly in regions with a small energy gap $\Delta_{10}(\mathbb{\Lambda})$. For drivings of isolated systems initiated in the ground state $|E_{0}(\mathbb{\Lambda}_{\rm I})\rangle$, the $v$=const. strategy improves the final fidelity for the target state $|E_{0}(\mathbb{\Lambda}_{\rm F})\rangle$ Matu23a ; Matu23b (see also Rola02 for a closely related method). We will show that this advantage lasts even at some nonzero (sufficiently low) temperatures. In principle, one can go beyond protocol B by using a curved path that minimizes the geometric length from $\mathbb{\Lambda}_{\rm I}$ to $\mathbb{\Lambda}_{\rm F}$, but this approach was shown to yield unsatisfactory results Matu23a and we do not apply it here.
@@
In Fig. 2 we plot the two driving paths in the plane $(\lambda,\chi)$ and the speeds (18) and (19) along both paths for driving protocols A and B for the system with ${N=10}$ qubits. We note the reduction of the planar speed $u(t)$ in both first- and second-order QPT regions for the $v$=const. driving procedure B. This slowdown is correlated with a reduced ground-state energy gap $\Delta_{10}$, the minimal gap for ${N=10}$ being however shifted with respect to the ${N\to\infty}$ QPT separatrix.